# Dynamics of inviscid annular supersonic jets

### v1

@@ Abstract
This article delves into the dynamics of inviscid annular supersonic jets, akin to those exiting converging-diverging nozzles in over-expanded regimes. It focuses on the first azimuthal Fourier mode of flow fluctuations and examines their behavior with varying mixing layer parameters and expansion regimes. The study reveals that two unstable Kelvin-Helmholtz waves exist in all cases, with the outer layer wave being more unstable due to velocity gradient differences. The inner layer wave is more sensitive to base flow changes and extends beyond the jet, potentially contributing to nozzle resonances. The article also investigates guided-jet modes, which are found to be robust and not highly sensitive to base flow parameters, making them essential for understanding jet dynamics. A simplified model is used to obtain ideal but base flows with realistic shape to study varying nozzle pressure ratios (NPR) effects on the dynamics of the waves supported by the jet.
@@ 1 Introduction
Resonances in jet flows originating from ”truncated ideal contour” (TIC) nozzles, and operating within the free separation regime, have been observed within very narrow over-expansion ratio ranges (Baars et al., 2012a; Jaunet et al., 2017; Martelli et al., 2020). Recent investigations have demonstrated that the associated pressure disturbances are likely to generate significant lateral forces that could lead to structural damage or even flight control problems (Bakulu et al., 2021). In these particular flow regimes, the presence of Mach disks in the flow decelerates the central core of the jet, while a higher-velocity annular flow envelops this slower core, as depicted in Figure 1. The mean flow thus comprises two co-annular jets, separated by a mixing layer emanating from the separation point and a slipline originating from the Mach disk triple point. The dynamics of these flow conditions have been shown to play a crucial role in the resonance mechanism and warrant more in-depth investigation to gain insights into the resonance process (Bakulu et al., 2021).
@@
Axisymmetric bulk velocity profile jets (single jets) have garnered significant attention within the scientific community, whether for their acoustics (Jordan & Colonius, 2013) or their dynamics. Substantial efforts have been dedicated to modeling the dynamics of these jets, employing linearized models, with the well-known $\tanh$ profile serving as a proxy for the jet’s base flow (Michalke, 1984). Supersonic jet flows have also been extensively studied, and numerous dynamical models have been developed to understand their dynamics (Tam & Hu, 1989), noise characteristics (Tam, 1972, 1991), and resonances (please refer to Edgington-Mitchell (2019) for a recent review).
Coaxial jets have also been a focal point for the scientific community, particularly during the late 1970s when it was discovered that they can exhibit reduced noise levels compared to their equivalent single jet counterparts under optimal operating conditions Dosanjh et al. (1969, 1971); Yu & Dosanjh (1971). Extensive experimental studies have shown that supersonic coaxial jets operated with a faster external stream, denoted as inverted velocity profile (IVP) jets, can lower the overall jet noise due to a decrease in shock-associated noise. To elucidate this phenomenon, a series of experimental and theoretical investigations ensued (Tanna et al., 1985; Tam & Tanna, 1985a, b).
@@
Given the similarities between the nozzle exhaust flow and IVP jets, it is reasonable to expect that the dynamics of nozzle jet flow may also exhibit resemblances to IVP jets. However, it’s worth noting that in Dahl & Morris (1997b), the study only considered cases where both flow streams were supersonic, while in the nozzle exhaust, the inner flow is decelerated to subsonic speed through a normal shock wave as depicted in figure reffig:flow. Additionally, recent studies have highlighted the importance of guided jet modes (GJM), neutral duct-like modes supported by the jet Towne et al. (2017); Schmidt et al. (2017), in the dynamics of supersonic jet flows (Gojon et al., 2018; Edgington-Mitchell et al., 2018; Jaunet et al., 2019; Mancinelli et al., 2021; Nogueira et al., 2022). These waves were first mentioned by Tam & Hu (1989), only two years before the study by Dahl & Morris (1997b), and were therefore not discussed in the latter paper.
To the best of our knowledge, no model has yet been proposed to predict the occurrence of resonances in the exhaust of an over-expanded convergent-divergent nozzle. Therefore, in this study, we propose to investigate the linear dynamics of over-expanded nozzle exhaust flows, with a particular focus on the effect of the base flow changes, mimicking variations in nozzle expansion regime, to the characteristics of the most unstable waves and guided jet modes (GJM).
@@ 2 Dynamical Model
2.1 Finite thickness model
We propose to use the same approach than Mancinelli et al. (2020), modelling the jet using locally-parallel linear stability theory. All variables are normalised by the fully expanded jet diameter $D_{j}$, the ambient density and speed of sound $\rho_{\infty}$ and $c_{\infty}$, respectively. The Reynolds decomposition,
$q^{\prime}(x,r,\theta,t)=\overline{q}(r)+q(x,r,\theta,t),$ (1)
$q(x,r,\theta,t)=\hat{q}(r)e^{i(kx+m\theta-\omega t)}$ (2)
where $k$ is the streamwise wavenumber and $m$ is the azimuthal Fourier wavenumber. The non-dimensional pulsation $\omega=2\pi StM_{a}$ is computed via the acoustic Mach number of the jet $M_{a}=U_{j}/c_{\infty}$, where $U_{j}$ is the equivalent fully expanded velocity of the flow, and the Strouhal number of the flow $St=f{D_{j}}/{U_{j}}$.
where $\gamma$ is the specific heat ratio for a perfect gas. The solution of the linear stability problem is obtained specifying a real or complex frequency $\omega$ and solving the resulting augmented eigenvalue problem $k=k(\omega)$, with $\hat{p}(r)$ the associated pressure eigenfunction. The eigenvalue problem is solved numerically by discretizing 3 in the radial direction using Chebyshev polynomials. A mapping function is used to non-uniformly distribute the grid points such that they are dense in the region of shear Trefethen (2000).
Although the model described above supports non-isobaric regime (Mancinelli et al., 2023), only isobaric jets will be studied in the following. Moreover, the article will only deal with fluctuations belonging the first helical mode ($m=1$), for this azimuthal mode is the only one responsible for side-loads in overexpanded nozzle flows (Dumnov, 1996), and for resonances observed in experiments (Jaunet et al., 2017).
2.2 Base flow parametrization
As can be seen in figure 2, the mean flow profiles downstream of the first Mach disk, are, as expected, composed of two MLs: one separating the supersonic flow and the ambient, the other one separating the subsonic core and the supersonic flow. A representative analytical formulation of the base flow is obtained through a regression of the numerical data available (DDES simulation previously done in Bakulu et al. (2021) with the following analytical profiles:
@@ 2.2 Base flow parametrization
where q here stands for $M$, $\bar{\rho}$ or $\bar{T}$. The indices i, a, and e respectively refer to the inner, the annular and the external flow variable of interest. $(R_{i},\theta_{qi})$,$(R_{e},\theta_{qe})$ the inner and external ML radial positions and thicknesses, respectively. As the flow velocity profile is concerned, this formulation is equivalent to the one used in Michalke (1984). This modelisation differs from the one chosen by Bhat & Seiner (1993) who preferred gaussian profiles instead of hyperbolic tangent ones.
Different mixing layer positions and thicknesses may be attributed the mean velocity, density or temperature profile when a direct fit to the numerical data is performed. This is shown in figure 2, where optimal analytical profiles are compared with the reference numerical data, demonstrating the ability of the analytical functions chosen to represent the computed flow field.
@@
2.3 Convergence study
We present in figure 3 the convergence of the eigenvalue spectrum with respect to the number of Chebychev collocation points $N$ and the Strouhal number. As can be seen, the convergence is more difficult to reach at high frequencies, especially for the Kelvin-Helmholtz modes (the isolated modes in the right lower quarter of the figure), but the spectrum seems converged above $N=301$ points. Therefore, a total number of $N=401$ collocation points is used for all the computations in this study.
Fig. 3: Convergence of the eigenvalue spectrum with respect to the number of Chebychev collocation points and Strouhal number from $0.1$ (blue) to $0.5$ (yellow).
3 Dynamics of annular supersonic jets
Contrary to a single stream jet flow, the base flow of a annular jet requires more parameters to be perfectly described. The characteristic features of the base flow (mixing layer thickness or position) may evolve with the nozzle pressure ratio as well and, therefore, their relative influence of the stability spectrum needs to be evaluated before one may explore the conditions for resonance. For this reason, we use the analytical profile described above and explore the influence of the main parameters on the stability properties of the jet.
3.1 Effect of radial position
The influence of the relative positions of the mixing layers is first studied by assuming a constant internal reference $u_{x}$ level. We perform linear stability analyses on base flows with varying inner mixing layer positions, as presented in Figure 4. The inner, annular, and external Mach numbers were chosen to approximately match the numerical simulations downstream of the first Mach disk: $M_{i}=0.6$, $M_{a}=1.4$, and $M_{e}=0$. The density profile was computed using the Crocco-Busemann relation (Michalke, 1984). A total of 36 base flows were used to finely study the changes in the corresponding linear dynamics.
Typical eigenspectra obtained at $St=0.2$ and $m=1$ are presented in Figure 5. On the right-hand side of the figure ($k_{r}>0$), two unstable eigenvalues corresponding to the Kelvin-Helmholtz (K-H) modes of both the inner and outer mixing layers can be recognized, as was found by Dahl & Morris (1997b) for supersonic IVP jets. The inner K-H mode is the least unstable of the two. The continuous acoustic branch is clearly visible, with a relatively weak influence of $R_{i}$ on the loci of the corresponding modes. Propagating and evanescent trapped wave wavenumbers (Towne et al., 2017) can be identified in the $k_{r}<0$ plane. Thus, the annular supersonic jets appear to exhibit similar features to those of a classical top-hat supersonic jet. In the following, we will focus on extracting the effect of the inner mixing layer position on the characteristics of these waves.
3.1.1 Kelvin-Helmholtz mode
In Figure 6, we plot the growth rate evolution of both K-H modes as a function of $R_{i}$ for various Strouhal numbers. The outer K-H mode is always more unstable than the inner one for all the configurations investigated in this study. This is logical considering that the velocity gradient is stronger for the outer mixing layer than for the inner one. The outer mode becomes even more unstable as $R_{i}$ increases, although at high frequencies its growth rate shows a plateau at low values of $R_{i}$. On the other hand, the inner mixing layer instability shows an optimal growth rate at specific $R_{i}$ values. This trend suggests that there may exist specific base flows and frequencies where the inner and outer K-H modes possess equivalent growth rates. In some cases, the inner mode can even become more unstable than the outer one. However, this is more likely to occur at higher frequencies and might not be relevant for the initial nozzle flow problem, where resonance at lower frequencies is observed.
The pressure eigenfunctions associated with the two identified K-H modes are presented in Figure 7 for $St=0.2$. As expected, the eigenfunction is maximum at the mixing layer location and is zero at the centerline, as we focus on the $m=1$ anti-symmetric wavenumber. The outer K-H mode does not seem to be significantly affected by the inner mixing layer, unlike the inner mode whose support spreads more through the outer mixing layer and above as the two mixing layers get closer. Interestingly, the radial support of the inner K-H wave decays less rapidly than that of the outer one. This allows the inner mode to be detected quite far from the jet and enables it to exchange energy with feedback waves outside of the supersonic annular region. This aligns well with the observations of Bakulu et al. (2021), who found that the inner mixing layer supports most of the perturbations associated with the resonance process in their flow.
3.1.2 Guided jet modes
Neutral waves, trapped by the jet (Towne et al., 2017), have been shown to play a significant role in the resonance of screeching and impinging jets (see, for example, Gojon et al. (2018); Jaunet et al. (2019); Mancinelli et al. (2021); Edgington-Mitchell et al. (2018)). Their dispersion relation is a key characteristic that explains cut-on and cut-off frequencies, the transition from axisymmetric to helical modes, and improves the model predictions (Mancinelli et al., 2019, 2021; Nogueira et al., 2022). Their eigenvalues lie on the real axis of the spectrum until they become evanescent. At supersonic speeds, jets support both upstream and downstream propagating modes. The upstream propagating mode carries energy in the upstream direction. Figure 5 shows that the annular jet configuration, as encountered in the exhaust of over-expanded CD nozzles, also supports such guided, trapped, neutral waves. The dispersion relation of these waves is plotted in Figure 8 in the $St-k_{r}$ plane for various positions of the inner mixing layer.
Fig. 8: Neutral mode dispersion relation for various positions of the inner mixing layer.
As expected, at each $R_{i}$, the neutral waves form families of modes hierarchically ordered by their radial supports: higher frequencies correspond to higher radial orders (i.e., the number of nodes and antinodes, see Tam & Hu (1989)). As seen, the position of the inner mixing layer has a strong impact on the neutral modes’ dispersion relation: as $R_{i}$ increases, the frequencies at which the modes are encountered also increase. Furthermore, the group velocity of the $k_{p}^{+}$ mode decreases with increasing $R_{i}$. This can be understood by recognizing that for $R_{i}=0$ or $R_{i}=R_{e}$, the base flow is close to a supersonic top-hat jet or a subsonic one, respectively. As can be seen, both group and phase velocities of the guided jet modes decreases with increasing $R_{i}$. This is logical if one considers the fact that increasing $R_{i}$ decreases the amount of high speed flow in the jet, hence, decreases the overall average speed of the jet. The neutral modes’ behavior seems to lie between what is expected for a subsonic $M=M_{i}<1.0$ jet and a supersonic $M=M_{a}>1.0$ one.
The eigenfunctions associated with the upstream propagating neutral modes are presented in Figure 9 for $St=0.1$ and $St=0.2$. As expected, the radial supports of these modes share common features with Bessel functions. At low Strouhal numbers, the eigenfunctions show one antinode, while for higher frequencies, more antinodes can be observed. Surprisingly, very little change in the radial support is observed for varying $R_{i}$. This is due to the fact that these modes are duct-like modes, so their support is mostly independent of the base flow, in contrast to their wavenumber. It is important to notice that they also have support outside the jet, similar to the inner K-H wave. This allows these waves to interact and for the upstream mode to carry energy upstream, potentially closing a feedback loop.
3.2 Effect of mixing layer thickness
In order to explore the influence of the thicknesses of the mixing layers on the stability properties of the flow, we vary the mixing layer thickness on a typical base flow as presented in Figure 10. Both mixing layers are centered around $D_{j}/2$ and $D_{j}/4$. The jet bulk velocities are chosen to be constant and comparable to those observed in the nozzle after the first Mach disk. This choice allows us to focus on the effect of the mixing layer thickness.
Fig. 10: Base flows used to study the effect of mixing layer thickness.
3.2.1 Kelvin-Helmholtz mode
Figure 11 shows the growth rate of both KH modes with varying thicknesses and Strouhal numbers. As expected, larger mixing layer thicknesses result in lower growth rates of the KH mode. For the base flows and frequencies considered here, the inner mixing layer is more stable than the outer one. This is due to their difference in velocity gradient. The outer mixing layer separates a supersonic stream from a quiescent atmosphere, while the inner one separates two rapid flows. Interestingly, for relatively high frequencies (above $St=0.2$), increasing the thickness stabilizes the outer mixing layer more quickly than the inner one. This is because the supersonic stream is compressed during the mixing layer thickening process. Thus, there might be flow configurations where both KH modes exhibit comparable growth rates. Depending on the frequency and base flow, the dynamics may not be entirely dominated by the inner KH mode, as one might initially assume.
The associated spatial profiles of the KH modes mentioned above are presented in Figure 12 for different Strouhal numbers and varying mixing layer thicknesses. We observe similar results to the previous study: for the thinnest mixing layer, the inner and outer KH modes exhibit maximum pressure fluctuations at the locations of their respective maximum velocity gradients. For all frequencies, thickening the mixing layer tends to spread the spatial support of the inner modes towards the outer mixing layer by increasing the pressure amplitude near the jet boundary. At the highest frequencies studied here, the signature of the outer KH modes also spreads onto the inner mixing layer. As indicated earlier by their relative growth rates, the entanglement of spatial support of the KH modes with increasing mixing layer thickness might make it difficult to distinguish their relative signatures in the fluctuating pressure field. Nonetheless, as mentioned in the previous section, the inner mixing layer KH mode appears to have a slower radial decay compared to the outer one. This supports the findings of Bakulu et al. (2021), which show that at the resonance frequency, the downstream energy-carrying mode was supported by the inner mixing layer. The wide radial extent of the eigenfunction, as shown here, also suggests that the inner KH mode can exchange energy with guided-jet modes or acoustic modes, whose support can be located outside the jet.
3.2.2 Guided jet modes
Similar to the radial position study of the mixing layer in the previous paragraph, the dispersion relation of the guided jet modes is presented in Figure 13. Unlike the previous cases, the thicknesses of the mixing layers have less effect on the dispersion relation of the guided jet modes. There is a slight variation in the domain of existence of the upstream-traveling modes between the branch and saddle points, as pointed out by Towne et al. (2017). The primary effect of thickening the mixing layers is on the group velocity ($\partial\omega/\partial k$), with thicker mixing layers leading to lower group velocities, as one would expect.
4 Evolution of dynamics with NPR
4.1 Modeling annular supersonic base flows
4.1.1 Derivation of the mean flow model
@@
Annular supersonic flows are commonly encountered when flows exit convergent-divergent nozzles and often exhibit a sufficiently large Mach disk within their core Hadjadj & Onofri (2009). Although references in the literature dedicated to analytically describe the topology of such shock wave network using mass, for example in Chow & Chang (1975); Li & Ben-Dor (1998), we opted to derive an even simpler description of the flow to obtain the parameter that our finite thickness model require. Indeed, the earlier-presented base flows can be parameterized as ($r_{i},r_{e},M_{i},M_{e},M_{a},T_{i},T_{a}$). Basic fluid mechanics principles reveal that these parameters are interconnected so we cannot assume arbitrary values. Hence, to attain a meaningful parameter set, a simplified model of the flow inside the nozzle is constructed, assuming quasi-one-dimensional behavior from the throat to the separation point, where the annular supersonic flow is presumed to originate. The simplified flow configuration, and the associated control volumes, that we propose to use in that purpose is illustrated in Figure 15.
@@
where $(\cdot)_{*}$ denotes sonic variables at the throat, and $D_{i,c}=\frac{D_{i}}{D_{e}}$ represents the normalized radius relative to the external vortex sheet. Note that $\mathcal{M}$ represents the local Mach number, differing from the acoustic Mach number employed in the linearized model: $\mathcal{M}=M\frac{c_{\infty}}{c}$, with $c$ signifying the local speed of sound. Consequently, the following relation can be easily obtained:
${\left(1+\frac{\gamma-1}{2}\right)^{-\frac{\gamma+1}{2(\gamma-1)}}}\frac{P_{t0}}{\sqrt{T_{t0}}}D_{c}^{2}$ $=$ $\frac{P_{i}}{\sqrt{T_{i}}}\mathcal{M}_{i}D_{i}^{2}+\frac{P_{a}}{\sqrt{T_{a}}}\mathcal{M}_{a}(1-D_{i}^{2})$ (6)
Assuming an isobaric jet for simplification i.e., $P_{a}=P_{i}=P_{\infty}$, the mass conservation equation can be further simplified:
@@
In essence, the mass conservation equation establishes a connection between the nozzle NPR, ${P_{t0}}/{P_{\infty}}$, its geometry and the Mach numbers of both the annular and internal jets. The NPR is dictated by the nozzle’s operational regime, while $D_{c}$ is given by nozzle design considerations.
@@
${\beta_{q}}D_{c}^{2}-\frac{T_{w}}{P_{\infty}}$ $=$ $\gamma D_{i}^{2}\left(\mathcal{M}_{i}^{2}-\mathcal{M}_{a}^{2}\right)+\left(\gamma\mathcal{M}_{a}^{2}+1\right),$ (8)
@@
and $T_{w}$ represents the pressure forces acting on the nozzle wall. It’s important to note that $T_{w}$ depends on the nozzle geometry and can be approximated using characteristic methods. By solving this equation, we can derive an analytical expression for $D_{i}^{2}$:
$D_{i}^{2}$ $=$ $\frac{1}{\gamma\left(\mathcal{M}_{i}^{2}-\mathcal{M}_{a}^{2}\right)}\left[{\beta_{q}}D_{c}^{2}-\frac{T_{w}}{P_{\infty}}-\left(\gamma\mathcal{M}_{a}^{2}+1\right)\right],$
which can subsequently be substituted into the mass conservation equation. This substitution leads to an equation for the annular Mach number:
${\frac{\mu_{i}-\mu_{a}}{\gamma\left(\mathcal{M}_{i}^{2}-\mathcal{M}_{a}^{2}\right)}\left[{\beta_{q}}{D_{c}^{2}}-\frac{T_{w}}{P_{\infty}}-\left(\gamma\mathcal{M}_{a}^{2}+1\right)\right]}+\mu_{a}-\beta_{m}$ $=$ $0,$ (10)
@@
It is worth mentioning that once $M_{i,a}$ are known, $T_{i,a}$ can be determined via the usual isentropic relations, using the typically available total temperature of the flow.
4.1.2 Solving the Model
Equation 10 can be solved to determine the Mach number of the annular flow, adhering to the fundamental principles of fluid mechanics. This can be achieved for any given input parameter set containing the nozzle pressure ratio ($NPR=\frac{P_{t0}}{P_{\infty}}$), throat diameter ($D_{c}$), inner subsonic Mach number ($\mathcal{M}_{i}$), and nozzle thrust ($T_{w}$). These parameters are intrinsically interconnected and must be chosen judiciously. We elucidate below how these parameters can be defined based on the nozzle’s pressure and Mach number profiles, along with the application of a classical separation criterion. Let us assume that flow separation transpires at an axial location within the nozzle, denoted as $x_{sep}$, where an isentropic expansion results in a local Mach number $\mathcal{M}_{sep}$ and pressure $P_{sep}$.
Primarily, it is evident from Figure 15 that our reference length scale, $D_{e}$, can be equated to the diameter of the nozzle wall at the separation point:

### v2

# Linear dynamics of over-expanded annular supersonic jets

@@ Abstract
This article delves into the dynamics of inviscid annular supersonic jets, akin to those exiting converging-diverging nozzles in over-expanded regimes. It focuses on the first azimuthal Fourier mode of flow fluctuations and examines their behavior with varying mixing layer parameters and expansion regimes. The study reveals that two unstable Kelvin-Helmholtz waves exist in all cases, with the outer layer wave being more unstable due to velocity gradient differences. The inner layer wave is more sensitive to base flow changes and extends beyond the jet, potentially contributing to nozzle resonances. The article also investigates upstream propagating guided-jet modes, which are found to be robust and not highly sensitive to base flow changes, making them essential for understanding jet dynamics. A simplified model is used to obtain ideal but base flows with realistic shape to study varying nozzle pressure ratios (NPR) effects on the dynamics of the waves supported by the jet.
@@ 1 Introduction
Resonances in jet flows originating from ”truncated ideal contour” (TIC) nozzles, and operating within the free separation regime, have been observed within very narrow over-expansion ratio ranges (Baars et al., 2012a; Jaunet et al., 2017; Martelli et al., 2020). Recent investigations have demonstrated that the associated pressure disturbances are likely to generate significant lateral, off-axis, forces that could lead to structural damage or even flight control problems (Bakulu et al., 2021). Despite the research efforts, no consensus on the origin of these resonances have yet emerged and among the various possible explanations the role of vortex shed by the Mach disk (Martelli et al., 2020), the effect of the separated mixing layer impingement in the nozzle lip (Tarsia Morisco et al., 2023) or an internal screech feedback loop (Jaunet et al., 2017) were hypothesized. The seek for the necessary ingredient for a feedback loop to be sustained in these atypical flows is one of the motivations for this study.
In these particular flow regimes, the presence of Mach disks in the flow decelerates the central core of the jet, while a higher-velocity annular flow envelops this slower core, as depicted in figure 1. The mean flow thus comprises two co-annular jets, separated by a mixing layer emanating from the separation point and a slipline originating from the Mach disk triple point. The dynamics of these flow conditions have been shown to play a crucial role in the resonance mechanism and warrant more in-depth investigation to gain insights into the resonance process (Bakulu et al., 2021).
@@
Axisymmetric bulk velocity profile jets (single jets) have garnered significant attention within the scientific community, especially to predict their acoustics (Jordan & Colonius, 2013). Hence, substantial efforts have been dedicated to modeling the dynamics of these jets, employing linearized models, with the well-known $\tanh$ profile serving as a proxy for the jet’s base flow (Michalke, 1984). Supersonic jet flows have also been extensively studied, and numerous dynamical models have been developed to understand their dynamics (Tam & Hu, 1989), noise characteristics (Tam, 1972, 1991), and resonances (please refer to Edgington-Mitchell (2019) for a recent review).
Coaxial jets have also been a focal point for the scientific community, particularly during the late 1970s when it was discovered that they can exhibit reduced noise levels compared to their equivalent single jet counterparts under optimal operating conditions (Dosanjh et al., 1969, 1971; Yu & Dosanjh, 1971). Extensive experimental studies have shown that supersonic coaxial jets operated with a faster external stream, denoted as inverted velocity profile (IVP) jets, can lower the overall jet noise due to a decrease in shock-associated noise. To elucidate this phenomenon, a series of experimental and theoretical investigations ensued (Tanna et al., 1985; Tam & Tanna, 1985a, b).
@@
Given the similarities between the nozzle exhaust flow and IVP jets, it is reasonable to expect that the dynamics of nozzle jet flow may also exhibit resemblances to the IVP jets one. However, it’s worth noting that in Dahl & Morris (1997b), the study only considered cases where both flow streams were supersonic, while in the nozzle exhaust, the inner flow is decelerated to subsonic speed through a normal shock wave as depicted in figure 1. Additionally, recent studies have highlighted the importance of guided jet modes (GJM) in the dynamics of supersonic jet flows. These waves were first mentioned by Tam & Hu (1989), only two years before the study by Dahl & Morris (1997b), and were therefore not discussed in the latter paper. Contrary the KH waves, these instability waves are neutral duct-like modes (Towne et al., 2017; Schmidt et al., 2017) that can have the interesting property of propagating in the upstream direction, towards the nozzle. This makes them perfect candidates to provide feedback in a resonance loop (Gojon et al., 2018; Edgington-Mitchell et al., 2018; Jaunet et al., 2019; Mancinelli et al., 2021; Nogueira et al., 2022).
To the best of the authors knowledge, there is no description of such waves in supersonic over-expanded jet, and no model has yet been proposed to predict the occurrence of resonances in the exhaust flow of convergent-divergent nozzle. Therefore, in this study, we propose to investigate the linear dynamics of over-expanded nozzle exhaust flows. A simplified framework of parallel flows is adopted and the study focuses on the effect of the base flow changes, mimicking variations in nozzle expansion regime, on the characteristics of the most unstable waves and the guided jet modes (GJM), which are essential components of some possible feedback resonance process.
@@ 2 Dynamical Model
2.1 Base flow modeling
2.1.1 Parametrization of the base flow
As can be seen in figure 2, the mean flow profiles downstream of the first Mach disk, are, as expected, composed of two MLs: one separating the supersonic flow and the ambient one, the other one separating the subsonic core and the supersonic flow. A representative analytical formulation of this base flow is obtained by connecting three uniform regions: the inner, the annular and the external one; with two hyperbolic tangent profiles to account for the mixing layers:
@@ 2.1.1 Parametrization of the base flow
where $q$ here stands for $M$, $\bar{\rho}$ or $\bar{T}$. The indices i, a, and e respectively refer to the inner, the annular and the external flow variable of interest. $(R_{i},\theta_{qi})$,$(R_{e},\theta_{qe})$ are the inner and external ML radial positions and thicknesses, respectively. A regression of the analytical profiles onto numerical data available (DDES simulation previously done in Bakulu et al. (2021)) gives access to the parameters. The reader should note that different mixing layer positions and thicknesses may be attributed to the mean velocity, density or temperature profile when a direct fit to the numerical data is performed. As the flow velocity profile is concerned, this formulation is equivalent to the one used in Michalke (1984). This modelisation differs from the one chosen by Bhat & Seiner (1993) who preferred gaussian profiles instead of hyperbolic tangent ones.
Optimal analytical profiles are compared with the reference numerical data in figure 2, demonstrating the ability of the analytical functions chosen to represent the computed flow field. The reader may note that the velocity and temperature profiles are not constant in the supersonic stream. This cannot be accounted for using the chosen analytical formulation and the modeled profiles show a plateau in this region. It is believed that this simplification will not lead to drastic change in the results as the velocity and temperature variation within the supersonic stream is very small compared to the variations in the mixing layers.
@@
2.1.2 Obtaining the base flow parameters
@@
Annular supersonic flows are commonly encountered when flows exit convergent-divergent nozzles and often exhibit a sufficiently large Mach disk within their core (Hadjadj & Onofri, 2009). A simplified analytical description of the shock waves network is proposed in various studies, using mass and momentum conservation, for example in for example in Chow & Chang (1975) and Li & Ben-Dor (1998), we opted to derive an even simpler description of the flow, using quasi 1D relations, to obtain the parameters that our finite thickness model requires. Indeed, the earlier-presented base flows can be parameterized with the following set of parameters: ($r_{i},r_{e},M_{i},M_{e},M_{a},T_{i},T_{a}$). In this section, we will often use comas to indicate interchangeable indices for sake conciseness. For example, the former vector of parameters could have been written ($r_{i,e},M_{i,e,a},T_{i,a}$) with this notation. Basic fluid mechanics principles reveal that these parameters are interconnected and we cannot assume arbitrary values. Hence, to attain a meaningful parameter set, a simplified model of the flow inside the nozzle is built, assuming quasi-one-dimensional behavior from the throat to the separation point, where the annular supersonic flow is presumed to originate. The simplified flow configuration, and the associated control volumes, that we propose to use in that purpose is illustrated in Figure 3.
@@
where $(\cdot)_{*}$ denotes sonic variables at the throat, and $D_{i,c}=\frac{D_{i,c}}{D_{e}}$ represents the normalized radius of the inner ($\cdot_{i}$) vortex sheet or the throat ($\cdot_{c}$) relative to the external vortex sheet.
Noting $\mathcal{M}$, the local Mach number, differing from the acoustic Mach number employed in the linearized model: $\mathcal{M}=M\frac{c_{\infty}}{c}$, where $c$ is the local speed of sound, the following relation can be easily obtained from the previous equation:
${\left(1+\frac{\gamma-1}{2}\right)^{-\frac{\gamma+1}{2(\gamma-1)}}}\frac{P_{t0}}{\sqrt{T_{t0}}}D_{c}^{2}$ $=$ $\frac{P_{i}}{\sqrt{T_{i}}}\mathcal{M}_{i}D_{i}^{2}+\frac{P_{a}}{\sqrt{T_{a}}}\mathcal{M}_{a}(1-D_{i}^{2}),$ (3)
where $P_{t0}$ and $T_{t0}$ are the total pressure and temperature of the flow and $P_{i,a}$ and $T_{i,a}$ are the static pressure and temperature of inner ($\cdot_{,}i$) and annular ($\cdot{,a}$) streams. Assuming an isobaric jet for simplification i.e., $P_{a}=P_{i}=P_{\infty}$, the mass conservation equation can be further simplified (see the appendix Appendix for further details):
@@
In essence, the mass conservation equation establishes a connection between the nozzle pressure ratio $NPR={P_{t0}}/{P_{\infty}}$, its geometry and the Mach numbers of both the annular and internal jets. The NPR is dictated by the nozzle’s operational regime, $D_{c}$ is given by nozzle design considerations.
@@
${\beta_{q}}D_{c}^{2}-\frac{F_{w}}{P_{\infty}}$ $=$ $\gamma D_{i}^{2}\left(\mathcal{M}_{i}^{2}-\mathcal{M}_{a}^{2}\right)+\left(\gamma\mathcal{M}_{a}^{2}+1\right),$ (5)
@@
and $F_{w}$ represents the pressure forces acting on the nozzle wall. It’s important to note that $F_{w}$ depends on the nozzle geometry and can be approximated using characteristic methods. By solving this equation, we can derive an analytical expression for $D_{i}^{2}$:
$D_{i}^{2}$ $=$ $\frac{1}{\gamma\left(\mathcal{M}_{i}^{2}-\mathcal{M}_{a}^{2}\right)}\left[{\beta_{q}}D_{c}^{2}-\frac{F_{w}}{P_{\infty}}-\left(\gamma\mathcal{M}_{a}^{2}+1\right)\right],$
which can subsequently be substituted into the mass conservation equation. This substitution leads to an equation for the annular Mach number, providing that $\mathcal{M}_{i}$ and $F_{w}$ are known:
${\frac{\mu_{i}-\mu_{a}}{\gamma\left(\mathcal{M}_{i}^{2}-\mathcal{M}_{a}^{2}\right)}\left[{\beta_{q}}{D_{c}^{2}}-\frac{F_{w}}{P_{\infty}}-\left(\gamma\mathcal{M}_{a}^{2}+1\right)\right]}+\mu_{a}-\beta_{m}$ $=$ $0,$ (7)
@@
It is worth mentioning that once $M_{i,a}$ are known, $T_{i,a}$ can be determined via the usual isentropic relations, using the typically known total temperature of the flow.
2.1.3 Solving the base flow model
Equation 7 can be solved to determine the Mach number of the annular flow, adhering to the fundamental principles of fluid mechanics. This can be achieved for any given input parameter set containing the nozzle pressure ratio ($NPR=\frac{P_{t0}}{P_{\infty}}$), throat diameter ($D_{c}$), inner subsonic Mach number ($\mathcal{M}_{i}$), and nozzle thrust ($F_{w}$). These parameters are intrinsically interconnected and must be chosen judiciously. We elucidate below how these parameters can be defined based on the nozzle’s pressure and Mach number profiles, along with the application of a classical separation criterion. Let us assume that flow separation occurs at an axial location within the nozzle, denoted as $x_{sep}$, where an isentropic expansion of the flow inside the nozzle has resulted in a local Mach number $\mathcal{M}_{sep}$ and pressure $P_{sep}$.
Primarily, it is evident from Figure 3 that our reference length scale, $D_{e}$, can be equated to the diameter of the nozzle at the separation point $D_{w}$: